# note-name = , use-sort-key = false

### v3

@@ Appendix B Deriving the system’s Wigner function equation of motion
We refer mainly to Refs. [43, 85] for the formalism in this section. The equation of motion for a Wigner function can be obtained by applying the Wigner transformation to the Hilbert-space equation of motion. The Wigner transformation is linear—the Wigner transform of a sum is a sum of the Wigner transforms of the summands. What about the Wigner transform of a product? The answer is the Moyal $\star$-product of the Wigner transforms of the factors. The Moyal $\star$-product operator is defined as [43]
$\star=\exp\left[\frac{i}{2}\left(\overset{\leftarrow}{\partial_{x}}\overset{\rightarrow}{\partial_{p}}-\overset{\leftarrow}{\partial_{p}}\overset{\rightarrow}{\partial_{x}}\right)\right]$ (32)
where the overset arrows dictate the operation direction. The $\star$-product essentially multiplies two scalars while behaving like matrix multiplication—it is noncommutative, associative, linear, etc. To take the Wigner transform of some function $f\left(\hat{A}_{0},\hat{A}_{1},\dots\right)$ of Hilbert space operators, we simply replace the operators by their Wigner transforms $W_{\hat{A}_{0}},W_{\hat{A}_{1}},\dots$ and replace all product operations with $\star$-products to obtain the $\star$-function” $f_{\star}\left(W_{\hat{A}_{0}},W_{\hat{A}_{1}},\dots\right)$. Evaluating the $\star$-product using its definition may be cumbersome. One technique is to utilize the “Bopp shift”. Based on the property that
$\exp\left(\hat{A}\partial_{x}\right)f(x)=f(x+\hat{A})$ (33)
@@
$\begin{split}f(x,p)\star g(x,p)&=f\left(x+\frac{i}{2}\overset{\rightarrow}{\partial}_{p},p-\frac{i}{2}\overset{\rightarrow}{\partial}_{x}\right)g(x,p)\\ &=f\left(x+\frac{i}{2}\overset{\rightarrow}{\partial}_{p},p\right)g\left(x-\frac{i}{2}\overset{\leftarrow}{\partial}_{p},p\right)\\ &=f\left(x,p-\frac{i}{2}\overset{\rightarrow}{\partial}_{x}\right)g\left(x,p+\frac{i}{2}\overset{\leftarrow}{\partial}_{x}\right)\\ &=f(x,p)g\left(x-\frac{i}{2}\overset{\leftarrow}{\partial}_{p},p+\frac{i}{2}\overset{\leftarrow}{\partial}_{x}\right)\end{split}$ (34)
@@
$\begin{split}\rho^{\prime}&=-i\left[\hat{H},\rho\right]+\sum_{k}\gamma_{k}\left(\hat{O}_{k}\rho\hat{O}_{k}^{\dagger}-\frac{1}{2}\left\{{\hat{O}_{k}^{\dagger}\hat{O}_{k}},{\rho}\right\}\right)\\ &\Bigg{\downarrow}\text{(Wigner transform)}\\ \partial_{t}W&=-i\left(W_{\hat{H}}\star W-W\star W_{\hat{H}}\right)\\ &\quad+\sum_{k}\gamma_{k}\Bigg{(}W_{\hat{O}_{k}}\star W\star W_{\hat{O}_{k}}^{*}\\ &\qquad\qquad\qquad-\frac{1}{2}W_{\hat{O}_{k}}^{*}\star W_{\hat{O}_{k}}\star W\\ &\qquad\qquad\qquad-\frac{1}{2}W\star W_{\hat{O}_{k}}^{*}\star W_{\hat{O}_{k}}\Bigg{)}\\ \end{split}$ (35)
where we have used the property that $W_{\hat{A}^{\dagger}}=W_{\hat{A}}^{*}$. This equation can be more concisely written using the Moyal bracket $\left\{\left\{{f},{g}\right\}\right\}=-i\left(f\star g-g\star f\right)$ which is the extension of the Poisson bracket to quantum mechanics. Furthermore, as the term $\left(f\star g-g\star f\right)$ is called the $\star$-commutation”, we may define the $\star$-dissipator” $\mathcal{D}_{\star}\left(\hat{O}\right)[\rho]$ to shorthand the Wigner transforms of the Lindblad dissipators. We may thus write
$\partial_{t}W=\left\{\left\{{W_{\hat{H}}},{W}\right\}\right\}+\sum_{k}\gamma_{k}\mathcal{D}_{\star}\left(\hat{O}_{k}\right)[\rho]$ (36)
@@
$W_{\hat{a}}$ $=\frac{1}{\sqrt{2}}\left(W_{\hat{x}}+iW_{\hat{p}}\right)=\frac{x+ip}{\sqrt{2}}$ (37a)
$W_{\hat{a}^{\dagger}}$ $=W_{\hat{a}}^{*}=\frac{x-ip}{\sqrt{2}}$ (37b)
$W_{\hat{a}^{\dagger}\hat{a}}$ $=W_{\hat{a}^{\dagger}}\star W_{\hat{a}}=\frac{x^{2}+p^{2}-1}{2}$ (37c)
$W_{\hat{a}\hat{a}^{\dagger}}$ $=W_{\hat{a}}\star W_{\hat{a}^{\dagger}}=\frac{x^{2}+p^{2}+1}{2}=1+W_{\hat{a}^{\dagger}}\star W_{\hat{a}}$ (37d)
$W_{\hat{a}^{2}}$ $=W_{\hat{a}}\star W_{\hat{a}}=\frac{x^{2}+2ixp-p^{2}}{2}=W_{\hat{a}}^{2}$ (37e)
$W_{\hat{a}^{\dagger{2}}\hat{a}^{2}}$ $=W_{\hat{a}^{\dagger{2}}}\star W_{\hat{a}^{2}}$
$=\frac{x^{4}+2p^{2}x^{2}+p^{4}-4x^{2}-4p^{2}+2}{4}$ (37f)
$=W_{\hat{a}^{\dagger}\hat{a}}\star W_{\hat{a}^{\dagger}\hat{a}}-W_{\hat{a}^{\dagger}\hat{a}}$
It is then straightforward—though lengthy—to evaluate the equation of motion for the system’s Wigner function using these equations. As an example, we have
$\begin{split}W_{\hat{a}^{\dagger}\hat{a}}\star W&=\frac{1}{2}\left[\left(x+\frac{i}{2}\overset{\rightarrow}{\partial}_{p}\right)^{2}+\left(p-\frac{i}{2}\overset{\rightarrow}{\partial}_{x}\right)^{2}-1\right]W\\ &=\left(\frac{x^{2}+p^{2}-1}{2}\right)W-\frac{i}{2}p\partial_{x}W+\frac{i}{2}x\partial_{p}W\\ &\quad-\frac{1}{8}\partial_{x}^{2}W-\frac{1}{8}\partial_{p}^{2}W\end{split}$ (38)
@@
$\begin{split}W\star W_{\hat{a}^{\dagger}\hat{a}}&=\frac{W}{2}\left[\left(x-\frac{i}{2}\overset{\leftarrow}{\partial}_{p}\right)^{2}+\left(p+\frac{i}{2}\overset{\leftarrow}{\partial}_{x}\right)^{2}-1\right]\\ &=\left(\frac{x^{2}+p^{2}-1}{2}\right)W+\frac{i}{2}p\partial_{x}W-\frac{i}{2}x\partial_{p}W\\ &\quad-\frac{1}{8}\partial_{x}^{2}W-\frac{1}{8}\partial_{p}^{2}W\end{split}$ (39)
@@
$\left\{\left\{{W_{\hat{H}=\hat{a}^{\dagger}\hat{a}+1/2}},{W}\right\}\right\}=-p\partial_{x}W+x\partial_{p}W$ (40)
Instead of evaluating the entire equation by hand, we utilize the SymPy module in Python. Our in-house package for the $\star$-product evaluation is available at https://github.com/hendry24/moyalstar.git. After post-programming algebraic manipulation, we obtain Eq. (15).

### v5

@@ Appendix B Deriving the system’s Wigner function equation of motion
We refer mainly to Refs. [44, 94] for the formalism in this section. The equation of motion for a Wigner function can be obtained by applying the Wigner transformation to the Hilbert-space equation of motion. The Wigner transformation is linear: the Wigner transform of a sum is a sum of the Wigner transforms of the summands. What about the Wigner transform of a product? The answer is the Moyal $\star$-product of the Wigner transforms of the factors. The Moyal $\star$-product operator is defined as [44]
$\star=\exp\left[\frac{i}{2}\left(\overset{\leftarrow}{\partial_{x}}\overset{\rightarrow}{\partial_{p}}-\overset{\leftarrow}{\partial_{p}}\overset{\rightarrow}{\partial_{x}}\right)\right],$ (36)
where the over-set arrows dictate the operation direction. The $\star$-product essentially multiplies two scalars while behaving like matrix multiplication: noncommutative, associative, linear, etc. To take the Wigner transform of some function $f\left(\hat{A}_{0},\hat{A}_{1},\dots\right)$ of Hilbert space operators, we simply replace the operators by their Wigner transforms $W_{\hat{A}_{0}},W_{\hat{A}_{1}},\dots$ and replace all product operations with $\star$-products to obtain the $\star$-function” $f_{\star}\left(W_{\hat{A}_{0}},W_{\hat{A}_{1}},\dots\right)$. Evaluating the $\star$-product using its definition may be cumbersome. One technique is to utilize the Bopp shift. Based on the property that
@@
given that $\hat{A}$ is independent of $x$, we may write  [^5] To see this more easily, consider the equivalent expression $f(x,p)\star g(x,p)=\exp\left[\frac{i}{2}\left(\partial_{x}\partial_{p^{\prime}}-\partial_{p}\partial_{x^{\prime}}\right)\right]f(x,p)g(x^{\prime},p^{\prime})|_{x^{\prime}=x,p^{\prime}=p}$.
$\begin{split}f(x,p)\star g(x,p)&=f\left(x+\frac{i}{2}\overset{\rightarrow}{\partial}_{p},p-\frac{i}{2}\overset{\rightarrow}{\partial}_{x}\right)g(x,p)\\ &=f\left(x+\frac{i}{2}\overset{\rightarrow}{\partial}_{p},p\right)g\left(x-\frac{i}{2}\overset{\leftarrow}{\partial}_{p},p\right)\\ &=f\left(x,p-\frac{i}{2}\overset{\rightarrow}{\partial}_{x}\right)g\left(x,p+\frac{i}{2}\overset{\leftarrow}{\partial}_{x}\right)\\ &=f(x,p)g\left(x-\frac{i}{2}\overset{\leftarrow}{\partial}_{p},p+\frac{i}{2}\overset{\leftarrow}{\partial}_{x}\right),\end{split}$ (38)
@@
[…]
where we have used the property that $W_{\hat{A}^{\dagger}}=W_{\hat{A}}^{*}$. This equation can be more concisely written using the Moyal bracket
$\left\{\left\{{f},{g}\right\}\right\}=-i\left(f\star g-g\star f\right),$ (40)
which is the extension of the Poisson bracket to quantum mechanics. Furthermore, as the term $\left(f\star g-g\star f\right)$ is called the $\star$-commutator”, we may analogously define the $\star$-dissipator” $\mathcal{D}_{\star}\left(\hat{O}\right)[\rho]$ to shorthand the Wigner transforms of the Lindblad dissipators. We may thus write
$\partial_{t}W=\left\{\left\{{W_{\hat{H}}},{W}\right\}\right\}+\sum_{k}\gamma_{k}\mathcal{D}_{\star}\left(\hat{O}_{k}\right)[\rho].$ (41)
@@
$W_{\hat{a}}$ $=\frac{1}{\sqrt{2}}\left(W_{\hat{x}}+iW_{\hat{p}}\right)=\frac{x+ip}{\sqrt{2}},$ (42a)
$W_{\hat{a}^{\dagger}}$ $=W_{\hat{a}}^{*}=\frac{x-ip}{\sqrt{2}},$ (42b)
$W_{\hat{a}^{\dagger}\hat{a}}$ $=W_{\hat{a}^{\dagger}}\star W_{\hat{a}}=\frac{x^{2}+p^{2}-1}{2},$ (42c)
and so on. It is then straightforward, though lengthy, to evaluate the equation of motion for the system’s Wigner function using these equations. As an example, we have
$\begin{split}W_{\hat{a}^{\dagger}\hat{a}}\star W&=\left(\frac{x^{2}+p^{2}-1}{2}\right)W-\frac{i}{2}p\partial_{x}W+\frac{i}{2}x\partial_{p}W\\ &\quad-\frac{1}{8}\partial_{x}^{2}W-\frac{1}{8}\partial_{p}^{2}W,\end{split}$ (43)
@@
$\begin{split}W\star W_{\hat{a}^{\dagger}\hat{a}}&=\left(\frac{x^{2}+p^{2}-1}{2}\right)W+\frac{i}{2}p\partial_{x}W-\frac{i}{2}x\partial_{p}W\\ &\quad-\frac{1}{8}\partial_{x}^{2}W-\frac{1}{8}\partial_{p}^{2}W,\end{split}$ (44)
@@
$\left\{\left\{{W_{\hat{H}=\hat{a}^{\dagger}\hat{a}+1/2}},{W}\right\}\right\}=-p\partial_{x}W+x\partial_{p}W.$ (45)
Instead of evaluating the entire equation by hand, we use our home-built symbolic package moyalstar [93], which utilizes the SymPy module in Python. After post-programming algebraic manipulation, we obtain Eq. (15).